Following are the data for the people on the Titanic listed by whether they survived and what type of ticket they had. Does survival appear to be independent of ticket class? Write the hypotheses and test it with an appropriate hypothesis test at 5% alpha level

First

Second

Third

Alive

202

118

178

Dead

123

167

528
Answer #1

H0: Survival appear to be independent of ticket class

vs

H1: Survival appear to be dependent of ticket class

Chi square formula

Expected frequencies = (Row total* Column Total) /Grand Total.

Degree of freedom =(Row-1)(Column-1) =(2-1)*(3-1)=2

Chi square test statistic value is =131.08

P value is approximately 0 .............................by using CHIDIST(131.08,2)

P value =0 < 0.05

Therefore, we reject H0 at

Conclusion: We can say that at Survival appear to be dependent of ticket class